OpenStudy (anonymous):

hi can someone pls. explain to me about Inverse trigonometric functions(Differential Calculus). pls. solve this problem then explain pls. y=square root of Arcsinx y=xArcsin2x

OpenStudy (anonymous):

now clear what the problem is so "solve" are you looking for the derivatives of these?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that is wrong. \[\frac{d}{dx}\sqrt{f(x)}=\frac{f'(x)}{2\sqrt{f(x)}}\]

OpenStudy (anonymous):

in this case \[f(x)=\arcsin(x)\] so \[f'(x)=\frac{1}{\sqrt{1-x^2}}\]

OpenStudy (anonymous):

so the first step is to write \[\frac{d}{dx}\sqrt{\arcsin(x)}=\frac{\frac{1}{\sqrt{1-x^2}}}{2\sqrt{\arcsin(x)}}\]

OpenStudy (anonymous):

then some algebra maybe to clean it up

OpenStudy (anonymous):

for \[y=x\arcsin(2x)\] you need the product rule and the chain rule

OpenStudy (anonymous):

product rule use \[(fg)'=f'g+g'f\] here \[f(x)=x,f'(x)=1,g(x)=\arcsin(x),g'(x)=\frac{2}{\sqrt{1-4x^2}}\]

OpenStudy (anonymous):

put it together in the product rule to get the answer

OpenStudy (anonymous):

tnx satellite